North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size

Q6 - 29/30 March 2022

As I was only too well aware, Tuesday was a dismal grey day at Grosmont after last week's Spring sunshine, with thick cloud on the tops of the moor road from Pickering. The enthusiastic, full of energy, team that day was Bill Dobson, Trevor Wilford, Gordon Wells, Chris Lawson (yes, I was late again!), and Ian Pearson.

Bill and Ian continued where they left off last Thursday with the Q6, removing the old sleeve from the copper pipework and getting a new sleeve made, which incidentally was made from an old LNER boiler plug found under the workshop bench. Gordon had arrived just in time, so he got the job of brazing the new sleeve to the copper pipes. The pipe was then fitted up into the cab, joining the section from the manifold to the pipe under the cab and completing the steam supply pipe to the atomiser. Bill, Gordon and Trevor continued all day fitting and adjusting the clamps around the copper and steam heat pipes to clear the brake rods. In the cab of the Q6, Ian changed the lefthand gauge frame top klinger packing, then disappeared under the loco and oiled the eccentrics. He also tightened up the inner and outer steam cataract reverser gland nuts, as there was evidence of steam leakage when last in steam.

Chris who arrived late morning, cleaned the two left hand display cases on the north wall of Deviation shed before lunch. The other four will be dealt with in future weeks as part of improving the appearance of the building for visitors. Then, as it remained dry, he started renovation work on the large Deviation Shed name board on the running line side of the building, removing some of the flaking orange paint from its surround and giving the top part of the surround a coat of white primer. More work to be done next week if it remains dry.

The weather was even worse on Wednesday - a wet cold day with rain, hail and snow. Very Spring like! The enthusiasts that day were Bill Dobson, Steve Hyman, Ed Bolam, David Potter, Nigel Hall (now recovered from Covid), Jon Bradley and Ian Pearson.

Jon assisted Bill to pressurise the Q6 atomiser pipe system. After a lengthy time to get the pump working up to about 240lbs, they managed to get some results they didn't want - a few leaking joints. The main one was the cone fitting to the manifold valve which had to be disconnected and and blued in. This found several high spots on a machined surface. The surface was ground and made to fit. More of this work to do on other joints next week however.

NYMR services are due to start on 4 April.